

********************************************************************
UARS WEEKLY STATUS REPORT - 18 DEC 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
Reprocessing of all SUSIM experiment days to version 20 continues.
***********
PEM
***********
PEM is currently conducting tests of the OBC activation. PEM activation
occurs during daylight only, when there is adequate power generated from
the UARS solar array. Currently, the PEM magnetometer and medium energy
spectrometers (particles below 30 keV) activate correctly under OBC control.
Tests conducted on the high energy particle spectrometers (HEPS -- particles
greater than 30 keV) have had limited success. New OBC activation procedure
for HEPS based on previous tests has been created. Tests of the new activation
sequence will begin shortly. Due to testing, HEPS science data between
December 1 and December 15 should be considered invalid and not used.
***********
HALOE
***********
HALOE was powered on December 17, 1998. All engineering parameters are
nominal and science channel exoatmospheric values are within their expected
range. Science data processed results will be available the beginning of
next week.
HALOE science data version 19 reprocessing is complete. The data will be
available on the HALOE web site in January after the data validation is
complete.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
HRDI continues to operate normally, currently running in the 'daytime
only' mode.
Many thanks to all the Goddard folks who helped us during our recent
communications problems.
Since this is likely to be our last status report for 1998, we would like
to wish everyone Merry Christmas and a Happy New Year, and
continued funding in 1999.
***********
WINDII
***********
Version 5 of the WINDII production software is still undergoing
testing, and has not yet been promoted to production; that
should happen early in 1999. The WINDII Team extends the best
wishes of the Season, and for the New Year to all members of
the UARS Community.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The UK Meteorological Office has resumed sending UKMO data to the CDHF after
computing problems were resolved. No backlog exists.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
9 days HAL (4.80) PROD (L2-L3) (89R-89C=0L)
62 days HAL (4.80) PROD (L3) (1874R-1538C=336L)
41 days SOL (4.62) PROD (L2-L3) (2313R-2239C=74L)
18 days SUS (7.00) PROD (L1) (43R-43C=0L)
63 days SUS (7.00) PROD (L3) (2540R-886C-4F=1650L)
3 days CLA (7.00) ENG (L1-L3) (24R-24C=0L)
The CDHF has started the installation of IDL software to version 5.1.
The CDHF met with UCSS developers to discuss and try to determine why the
wallclock times of the SOLSTICE V4.60 production reprocessing is steadily
increasing with the mission day number. The CDHF demonstrated the UCSS
Scheduler methodology concerning the SOLSTICE scheduling and setup.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 05 December 1998 (Orbit 39522, GMT Day 345) through 11
December 1998 (Orbit 39627, GMT Day 345).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 342/0047 39568 On
HALOE 343/2307 39597 Off
The beta angle ranged between 37.0 degrees to 9.0 degrees for this report
period. The beta angle is now decreasing toward a minimum angle of -1.0
degrees on 13 December 1998 (DOY 347).
PEM successfully tested its latest revision, a 37 command HEPS
microprocessor RAM patch with RTS one-second delay timing every four
commands. The ability to do extra long RTS loads will reduce the overhead
in the existing RTS's and allow workarounds. The test results are being
evaluated.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.4 degrees C. Additional battery performance data
for each battery on 04 December 1998 and 11 December 1998 are provided in
Appendix D.
The clock error ranged from 9.9 to -10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HALOE 02
% PEM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496936
Transponder B: 98/300 - 2287.498964
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/337 show the observatory data loss to be
21 hours, 40 minutes, 14.260 seconds (an increase of 7.488 seconds since
last report period). This is a 0.0342 percent data loss which equals a
99.9658 percent data capture for the mission.
The increase of 7.488 seconds occurred over five (5) days:
7 0.224 seconds on Day 338 due to data hit;
7 1.536 seconds on Day 339 due to hit in data (IP);
7 1.696 seconds on Day 341 due to data hit and hit in data (IP);
7 2.272 seconds on Day 342 due to hit in data (IP); and
7 1.760 seconds on Day 344 due to hit in data (IP).
Two (2) new versions of Relative Time Sequence 32 (RTS 32) were generated
this week to refine the PMON controlled operation of HEPS1 portion of the
PEM instrument.
Flight Dynamics Facility (FDF) had to re-generate this week long term (8-
week) planning aids due to a last minute change by FOT in the planned
maneuver time for the upcoming UARS Roll Offset maneuver. The FOT time
change delayed planning aid delivery that subsequently resulted in the FOT
being behind schedule in much of their long-term planning tasks (i.e. TDRS
scheduling, SSPP Sunplan generation, etc).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Preparation of 0.5 degrees (colds side down) roll procedure to be executed
in December 1998 and January 1999 was completed and tested using the UARS
Test and Training Simulator (UTTS) on 08 December 1998. The UTTS testing
proceeded without error. The procedure and plan are now ready to be used
when scheduled on 351/14:55 UTC (17 December 1998).
Development of the procedure for the clock change to include the 01 January
1999 leap second adjustment continues.
FOT plans to begin testing of the Orbit Adjust Maneuver on the UTTS
(simulator) in December 1998. Two versions will be tested; one with B/D
thrusters and; one with A/C thrusters. The next orbit adjust will not be
needed until late 2000 (or later if the altitude limits are lowered).
Flight Operations and SOLSTICE have completed and verified the plan for
SOLSTICE Day-Only Operation. SOLSTICE will select a date for transition to
day-only operation.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - A set of long-term PSAT and S/C Orbital
Events files for Y2K testing which have consistent orbit numbers has been
received. All previously delivered sets of PL and OL files for the 1999-
to-2000 century crossover had orbit numbers which did not line up with the
other files, making it impossible to process them on the CMS. Now that the
FOT has a correct set of these two files, they can proceed with Y2K testing
and hopefully finish up all of the test functions related to these
products.
Flight Dynamics Facility (FDF) - see CMS paragraph.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Testing of the Trending System v.1.1 release (CM-0009) continued.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 13
December 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 December 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39524 339/0228 98-486 OBC Clk Adj fr 0.00 to 20.6 msec/day
39532 339/1455 98-487 HALOE Update SR & SS due to CS 4 error
39548 340/1608 98-488 OBC Clk Adj fr 20.6 to 0.00 msec/day
39568 342/0047 98-489 WINDII Power ON
39577 342/1605 98-490 PEM PEM RTS loads
39595 343/1945 98-491 PEM PEM RTS loads
39597 343/2307 98-492 HALOE Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 December 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HACS4 Y-HI Control status word 4 39532,33
HRPSEQBIMON Y-HI PSE QB Current Monitor 39588
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 December 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39524 339/02:28 (05 Dec 98) from 00.0 to 20.6 msec/day
39548 340/16:08 (06 Dec 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 December 1998
04 Dec 1998 - GMT Day 338 Beta = 37.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +84.0/+56.0 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +28.0/-22.4 +5.47/+4.51 +29.6/-15.2 25.8 1.027 21.0
3 +56.0/-5.6 +4.51/+2.92 +28.8/-15.2 25.8 1.026 20.7
11 Dec 1998 - GMT Day 345 Beta = 9.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +78.4/+56.0 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-5.6 +5.47/+4.19 +31.6/-14.4 25.4 1.020 21.5
3 +134.4/0.0 +4.19/+2.92 +30.4/-14.4 25.4 1.020 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 December 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-121 FOT 39531 339/1243 N/A Halo CS4 error
98-117 FOT 39534 339/1812 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-120 FOT 39521 338/21:15 OP ERROR INST LD
98-121 FOT 39532 339/14:55 HALOE CS4 ERROR
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT DEC 98 See List 14170 GENERIC LATE ACQ
SEVNTFW50